OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 12 and the last term is –3,359,232?

OpenStudy (matt6288):

What is the sum of an 8–term geometric series if the first term is 12 and the last term is –3,359,232 nth term = a r ^ ( n - 1 ) 8th term = 12 r ^ ( 8 - 1 ) = –3,359,232 12 r ^ ( 7 ) = –3,359,232 r^7 = -279,936 r = – 6

OpenStudy (anonymous):

its not giving me answer when i plug it in

OpenStudy (matt6288):

n Σ a r^k = a ( 1 - r^(n+1)) / ( 1 - r ) k=0 7 Σ 12 r^k = 12 ( 1 - ( – 6 ) ^(7+1)) / ( 1 - ( – 6 )) k=0 ...............= 12 ( 1 - ( – 6 ) ^(8)) / 7 ...............= 12 ( 1 - ( 1,679,616 )) / 7 ...............= 12 ( –1,679,615 ) / 7 ...............= –20,155,380 / 7 ...............= –2,879,340